# Improved Belief Propagation Decoding Algorithms for Surface Codes

### v3

@@ III Belief Propagation Optimization
Surface codes violate the assumptions of BP’s objective function and probability independence, making BP decoding difficult to converge. These issues manifest in symmetric trapping sets, where the local structure of each node is identical, and message updates are symmetric under specific error patterns. Breaking the former symmetry involves modifying the code structure, while in this section, we focus on breaking the latter, the symmetry of message updates.
III.1 Message update optimization
III.1.1 Update smoothing
@@
where $x_{t}$ is the observation at time $t$, $v_{t}$ is the smoothed value, and $\gamma$ is a decay factor.
In machine learning, momentum methods utilize an idea similar to EWA, adding a momentum term to accumulate historical information in gradient descent
$m_{t}$ $=\gamma m_{t-1}+(1-\gamma)\nabla J(\theta),$ (14)
$\theta$ $=\theta-\alpha m_{t},$ (15)
where $\gamma$ is the decay factor, $m_{t}$ is the accumulated momentum term, $\nabla J(\theta)$ is the gradient of the loss function with respect to parameter $\theta$, and $\alpha$ is the learning rate. This method integrates previous directions while updating each parameter, thereby smoothing the update direction.
BP decoding can be analyzed from the perspective of the energy function. The high degeneracy of surface codes causes the energy function to have many equivalent global minima within a neighborhood. In the case of symmetric trapping sets, BP’s specific update step size causes the hard decision results to oscillate between adjacent erroneous solutions. We incorporate the concept of momentum into BP’s a posteriori probability updates to solve this issue by breaking the symmetry in BP’s message updates.
Momentum-BP can be implemented under both parallel and serial scheduling. Here, we present the latter in algorithm 1, where the initialization and check node update remain the same as in traditional BP. Our approach computes the a posteriori probability before performs the vertical update. The posterior update rule of Momentum-BP is shown as Eq. (16) and Eq. (17) where $g_{v_{j},e}^{(t)}$ are the accumulated momentum term and $Q_{v_{j},e}^{(t)}$ are the a posteriori probability for the $j$-th variable node regarding the Pauli error $e$ in the $t$-th iteration. Notice that we transformed Eq. (8), so that the accumulated term is the posterior value from the previous iteration. In the variable node update step, we subtract the original message sent from the check node rather than the momentum-smoothed message to ensure stability and prevent overcompensation.
Algorithm 1 Momentum-BP under Serial Scheduling over GF(4)
Input: Parity-check matrix $\boldsymbol{H}$ of $M\times N$, syndrome vector $\boldsymbol{s}=(1,-1)^{M}$, a priori probability vectors $\{p_{v_{j}}^{X},p_{v_{j}}^{Y},p_{v_{j}}^{Z}\}_{j=1}^{N}$, max iterations $iter_{max}$, learning rate $\alpha=1$, and momentum rate $\gamma\geq 0$.
@@
$g_{v_{j},e}^{(t)}$ $=\gamma g_{v_{j},e}^{(t-1)}+(1-\gamma)\frac{\partial J}{\partial Q_{v_{j},e}^{(t-1)}}$ (16)
$Q_{v_{j},e}^{(t)}$ $=Q_{v_{j},e}^{(t-1)}-\alpha g_{v_{j},e}^{(t)}$ (17)
@@
else repeat from serial update.
For the gradient calculation, we ignore the coefficient terms in Eq. (11), using the difference between the posterior updates calculated by traditional BP and the a posteriori probabilities from the previous iteration as the gradient for the current iteration, expanded in Eq. (19). When $\alpha=1$ and $\gamma=0$, Eq. (19) is equivalent to the traditional BP’s a posteriori update; when $10<\alpha<$ 1 and $\gamma=0$, it is equivalent to the EWA of a posteriori updates so far. In practice, to avoid the malice of hyper-parameter tuning and to enhance the interpretability of the formula, we can either fix $\alpha$ at 1 or $\gamma$ at 0.
$Q_{v_{j},e}^{(t)}$ $=Q_{v_{j},e}^{(t-1)}-\alpha\left(\gamma g_{v_{j},e}^{(t-1)}+(1-\gamma)(Q_{v_{j},e}^{(t-1)}-Q_{v_{j},e}^{(0)}-\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle e,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)})\right)$ (19)
@@
The properties of Momentum-BP can be explained from two perspectives. (1) Since it takes into account message updates from previous iterations, the method smooths the update direction in each iteration, preventing the corresponding energy function from oscillating between two points and gradually converging to the global minimum between them. (2) Node-wise momentum updates gradually break the symmetry of messages in the trapping sets shown in Fig. 2, allowing BP to escape from these trapping sets. The trapping set simulation in Section IV.1 will support those explanations.
III.1.2 Adaptive step size
@@
Oscillation of message updates in trapping sets can also be attributed to the step size. In traditional LLR-BP, all variable node posterior updates result from directly accumulating messages from neighboring check nodes, which can become symmetric and cause oscillations. From the gradient perspective, if the step size is too large, the optimization algorithm is likely to oscillate; if the step size is too small, the algorithm is likely unable to update effectively. In practical decoding, the update frequency of different variable nodes may be uneven, with some requiring more rapid updates while others need to remain stable.
@@
Inspired by adaptive gradient methods in deep learning, we propose AdaGrad-BP. The framework of this algorithm is very similar to Momentum-BP, with initialization and horizontal updates identical to traditional BP, while the posterior probability updates follow an adaptive step size rule, expressed as
$Q_{v_{j},e}^{(t)}=Q_{v_{j},e}^{(t-1)}-\alpha\frac{g_{v_{j},e}^{(t)}}{\sqrt{G_{v_{j},e}^{(t)}}+\epsilon},$ (21)
where $g_{v_{j},e}^{(t)}$ approximates the gradient of the energy function
@@
$g_{v_{j},e}^{(t)}=Q_{j,e}^{(t-1)}-Q_{j,e}^{(0)}-\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle e,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)},$ (22)
and $G_{v_{j},e}^{(t)}$ is the cumulative sum of the squared gradients as
$G_{v_{j},e}^{(t)}=G_{v_{j},e}^{(t-1)}+(g_{v_{j},e}^{(t)})^{2}.$ (23)
@@
This method adaptively adjusts the step size by accumulating historical gradient information. For instance, when the estimate of a variable node for a certain Pauli error changes significantly, the algorithm reduces the step size to counteract oscillation. Eq. (21) takes effect after the first iteration to ensure initial update stability. Here, $\alpha$ is the boost learning rate, providing an update impetus as the algorithm begins to accumulate gradients. This learning rate is relatively insensitive; a moderate value suffices. In all our tests, $\eta$ is fixed at 5. While AdaGrad-BP may face issues with accumulated gradients causing the step size to approach zero after many iterations, this does not significantly impact the algorithm. In practice, AdaGrad can reduce the number of iterations, allowing most error patterns to converge before the gradients become too small.
@@
Unfortunately, combining Momentum-BP and AdaGrad-BP using the Adam optimizer or other methods has not yielded satisfactory results. On one hand, the Adam algorithm introduces additional hyperparameters whose optimal values for BP decoding cannot be directly borrowed from deep learning, thereby complicating the algorithm. On the other hand, Momentum and AdaGrad approach similar message update optimizations from different perspectives, and a straightforward combination of the two can lead to overcompensation.
III.2 EWA Initialization of a priori probability
@@
BP decoding and gradient descent still have a fundamental difference, which limits the performance of algorithms in Section III.1. Each update in gradient descent is based on the previous step’s results, whereas in BP decoding, each calculation of a posteriori probabilities accumulates the same a priori probabilities. Given the objective function as shown in Eq. (2), for simplicity using the binary domain for instance, BP calculates the LLR values as
@@
$=Q_{v_{j}}^{(0)}+\sum_{\begin{subarray}{c}c_{i^{\prime}}\in\mathcal{M}(v_{j})\backslash c_{i}\end{subarray}}m_{c_{i^{\prime}}\rightarrow v_{j}},$ (26)
where $Q_{v_{j}}^{(0)}$ is a manually set initial value of the algorithm that participates in each iteration’s a posteriori probability update. However, since the prior error probabilities are derived from an theoretical channel which is hard to measure and estimate in experiments, the a priori probabilities can significantly affect the BP iterations.
Existing methods for altering initial probabilities include re-initialization using information obtained from previous iterations [22, 23, 31], as well as detecting oscillations during the decoding process and modifying initial values accordingly [32]. However, they either require multiple outer loops or manually set algorithm parameters. This subsection introduces a method for dynamically updating a priori probabilities, employing a transformed exponential weighted averaging, incorporating components of all previous a posteriori probabilities into the a priori probabilities with a decaying factor. Hence, we call this method the EWAInit-BP. EWAInit-BP under parallel scheduling is shown in algorithm 2.
Algorithm 2 EWAInit-BP under Parallel Scheduling over GF(4)
Input: Parity-check matrix $\boldsymbol{H}$, syndrome vector $\boldsymbol{s}$, a priori LLR vectors $Q^{(0)}$, max iterations $iter_{max}$, discount factor $\alpha$.
@@
EWA Initialization:
for $j=1\to N$, $e\in\{X,Y,Z\}$, do
$\Pi_{v_{j},e}^{(t)}$ $=\alpha Q_{v_{j},e}^{(0)}+(1-\alpha)Q_{v_{j},e}^{(t-1)}$ (27)
@@
for $j=1\to N$, calculate $m_{v_{j}\rightarrow c_{i}}^{(t)}$ as Eq. (7)
@@
for $j=1\to N$, $e\in\{X,Y,Z\}$, do
Calculate $Q_{v_{j},e}^{(t)}$ as Eq. (8)
if $Q_{v_{j},e}^{(t)}>0$ for all $e$, let $\hat{E}_{i}=I$
else, let $\hat{E}_{i}=\underset{e\in\{X,Y,Z\}}{\arg\min}Q_{v_{j},e}^{(t)}$
@@
else repeat from EWA Initialization.
In fact, combining Eq. (27) and Eq. (8), EWAInit-BP exhibits similar characteristics to Momentum-BP, namely the exponentially decaying components of previous iterations. The difference is that the former aggressively changes the ”prior” probability of each variable node in each iteration. Eq. (27) still retains a portion of the input prior probabilities. In our experiments, we tried replacing this term with the prior probabilities from the previous iteration, but the algorithm’s performance did not significantly change.
III.3 Complexity
@@ IV Simulation Results
IV.1 Breaking the Trapping Sets
@@
(b) Momentum-BP
(c) AdaGrad-BP
Figure 3: Comparison of traditional LLR-BP4, Momentum-BP, and AdaGrad-BP under parallel scheduling on the (4,0) trapping set. The latter two methods break the symmetry of message updates and converge within 12 iterations.
We first compare the performance of improved BP on the trapping set shown in Fig. 2. In this example, qubit 1 and qubit 2 experience X errors, resulting in an error syndrome of $[1,0,1,0]$. Fig. 3 shows the comparison between the traditional LLR-BP4, Momentum-BP and AdaGrad-BP in Section III.1 under parallel scheduling, with the maximum iteration set to 50. The hard decision result of the traditional BP oscillates continuously between all $X$ and all $I$, while improved BP algorithms show significantly reduced oscillation amplitude and converge within 12 iterations.
A more detailed comparison of the different algorithms proposed in this article is presented in Table. 1. The parameters used for these algorithms are defaults and were not specially optimized for this example. MBP [20] converges at the same fastest speed under serial scheduling but oscillates under parallel scheduling. BP-OTS [32] shows little dependency on scheduling method but its convergence depends on the manually set oscillation period $T$, which may vary in different error correction scenarios. Algorithms proposed in Section III.1, Momentum-BP and AdaGrad-BP, both converge under parallel scheduling. Besides, AdaGrad-BP and EWAInit-BP with adaptive a priori probabilities converge at the fastest speed under serial scheduling.
Table 1: Iterations required by different BP algorithms on the (4,0) trapping set
Algorithm  [^1] The parameters used for each algorithm are: MBP ($\alpha=0.6$), Momentum-BP ($\alpha=0.6$), AdaGrad-BP ($\alpha=5$), EWAInit-BP ($\alpha=0.6$) Parallel Scheduling Serial Scheduling
@@
MBP [20] 2
BP-OTS [32] T+2 T+2
EWAInit-BP 2
Momentum-BP 12 6
AdaGrad-BP 12 2
IV.2 Improved BP on Topological codes
In this subsection, we compare the performance of various BP improvements proposed in this article. Subsequently, we present comparisons of our algorithm with other BP improvements without post-processing for toric codes, planar surface codes, and XZZX surface codes. For all decoding algorithms, the maximum number of iterations $iter_{max}$ across all code lengths is set to 150 to ensure linear asymptotic complexity.
IV.2.1 Comparison of our BP optimization methods
Figure 4: Performance of various BP optimization methods presented in this article under serial scheduling, demonstrated using the $L=8$ toric code as an example. Each method uses default parameters. Among these, the method for optimizing initial values, EWAInit-BP, exhibits the best performance and completely eliminates the error floor.
A comparison of all methods proposed in this article is illustrated in Fig. 4. The two methods for optimizing message update rules, Momentum-BP and AdaGrad-BP, exhibit some improvements over traditional BP at low physical error rates, while their performance gradually converging to that of traditional BP as the physical error rate increases. The method for optimizing initial values, EWAInit-BP, demonstrate improvements ranging from 1 to 3 orders of magnitude over traditional BP, and excellently eliminate the error floor. However, the performance of algorithms that combine these two ideas is not satisfactory, because both approaches inherently consider the historical a posteriori probabilities but from different positions within the BP algorithm. This two methods are essentially addressing the same problem—incorporating past information to refine current updates. As a result, simply merging them can lead to redundancy and overcompensation, which negatively impacts the overall performance.
IV.2.2 Toric code
We construct the $[[2n^{2},1,n]]$ toric code based on its properties as a hypergraph product of an $[n,1]$ repetition code. The toric code features periodic boundaries, hence all its stabilizers have a weight of four, and any error syndromes on it exhibit translational symmetry [16], which is particularly conducive to BP and neural network decoding. We compare the performance of our EWAInit-BP and other existing methods under both parallel and serial scheduling as shown in Fig. 5 and Fig. 6, respectively.

### v5

@@ III Belief Propagation Optimization
From the perspective of BP as gradient descent [21], the high degeneracy of surface codes results in many equivalent global minima, leading to BP oscillating between adjacent erroneous solutions within symmetric trapping sets. A natural idea is to explore if certain empirical methods can be employed to smooth BP’s message updates, stabilizing the a posteriori probabilities at a specific local minimum before oscillations occur, thus enabling successful convergence to degenerate errors.
III-A Gradient Optimization of Message updates
This subsection follows the gradient perspective, inspired by optimizers in machine learning, to smooth BP iterations by modifying its message update rules.
III-A1 Update gradient smoothing
@@
where $x_{t}$ is the observation at time $t$, $v_{t}$ is the smoothed value, and $\gamma$ is a decay factor satisfying $10<\gamma<$ 1.
In machine learning, momentum methods utilize an idea similar to EWA, adding a momentum term to accumulate historical information in gradient descent as
$m_{t}$ $=\gamma m_{t-1}+(1-\gamma)\nabla J(\theta_{t-1}),$ (11)
$\theta_{t}$ $=\theta_{t-1}-\alpha m_{t},$ (12)
where $\gamma$ is the decay factor similar to that in EWA, $m_{t}$ is the accumulated momentum term, $\nabla J(\theta)$ is the gradient of the loss function with respect to parameter $\theta$, and $\alpha$ is the learning rate. This method integrates previous directions while updating parameters, thereby smoothing the update direction.
We incorporate the concept of momentum into BP’s a posteriori probability updates to solve this issue by breaking the symmetry in BP’s message updates, proposing Momentum-BP. Momentum-BP can be implemented under both parallel and serial scheduling. Here, we present the latter in Algorithm 1, where the initialization and check node update remain the same as in traditional BP. Our approach computes the a posteriori probability before performs the vertical update. The posterior update rule of Momentum-BP is shown as (13) and (14) where $g_{v_{j},W}^{(t)}$ are the accumulated momentum term and $Q_{v_{j},W}^{(t)}$ are the a posteriori probability for the $j$-th variable node regarding the Pauli error $W$ in the $t$-th iteration. Notice that we transformed (9), so that the accumulated term is the posterior value from the previous iteration. In the variable node update step, we subtract the original message sent from the check node rather than the momentum-smoothed message to ensure stability and prevent overcompensation.
For implementation of the gradient calculation, we ignore the coefficient terms in (29), using the difference between the posterior updates calculated by traditional BP and the a posteriori probabilities from the previous iteration as the gradient for the current iteration, expanded in (16). When $\alpha=1$ and $\gamma=0$, (16) is equivalent to the traditional BP’s a posteriori update; when $10<\alpha<$ 1 and $\gamma=0$, it is equivalent to the EWA of a posteriori updates. In practice, to avoid the malice of hyper-parameter tuning, we either fix $\alpha$ at 1 or $\gamma$ at 0.
The properties of Momentum-BP can be explained from two perspectives. (1) Since it takes into account message updates from previous iterations, the method smooths the update direction in each iteration, preventing the corresponding energy function from oscillating between two points and gradually converging to the global minimum. (2) Node-wise momentum updates gradually break the symmetry of message passing shown in Fig. 1, allowing BP to escape from these trapping sets. The trapping set simulation in Section IV-A will support those explanations.
Algorithm 1 Momentum-BP under Serial Scheduling
Input: Parity-check matrix $\boldsymbol{H}$ of $M\times N$, syndrome vector $\boldsymbol{s}=(1,-1)^{M}$, a priori LLR vectors $\{\Pi_{v_{j}}^{(0)}\}_{j=1}^{N}$, max iterations $iter_{max}$, learning rate $\alpha=1$, and momentum rate $\gamma\geq 0$.
@@
$g_{v_{j},W}^{(t)}$ $=\gamma g_{v_{j},W}^{(t-1)}+(1-\gamma)\frac{\partial J}{\partial Q_{v_{j},W}^{(t-1)}}$ (13)
$Q_{v_{j},W}^{(t)}$ $=Q_{v_{j},W}^{(t-1)}-\alpha g_{v_{j},W}^{(t)}$ (14)
@@
else $t=t+1$, repeat from serial update.
$Q_{v_{j},W}^{(t)}$ $=Q_{v_{j},W}^{(t-1)}-\alpha\left(\gamma g_{v_{j},W}^{(t-1)}+(1-\gamma)(Q_{v_{j},W}^{(t-1)}-\Pi_{v_{j},W}-\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)})\right)$ (16)
@@
[…]
III-A2 Adaptive step size
@@
Oscillation of message updates in trapping sets can also be attributed to the step size. In traditional LLR-BP, all variable node posterior updates result from simply directly accumulating messages from neighboring check nodes. From the gradient perspective, if the step size is too large, the optimization algorithm is likely to oscillate; if the step size is too small, the algorithm is likely unable to update effectively. In practical decoding, the update frequency of different variable nodes may be uneven, with some requiring more rapid updates while others need to remain stable.
@@
Inspired by adaptive gradient methods in deep learning, we propose AdaGrad-BP. The framework of this algorithm is similar to Momentum-BP, with initialization and horizontal updates identical to traditional BP, while the posterior probability updates follow an adaptive step size rule as
$Q_{v_{j},W}^{(t)}=Q_{v_{j},W}^{(t-1)}-\alpha\frac{{g^{{}^{\prime}}}_{v_{j},W}^{(t)}}{\sqrt{G_{v_{j},W}^{(t)}}+\epsilon},$ (19)
where ${g^{{}^{\prime}}}_{v_{j},W}^{(t)}$ approximates the gradient of energy function
@@
${g^{{}^{\prime}}}_{v_{j},W}^{(t)}=Q_{v_{j},W}^{(t-1)}-\Pi_{v_{j},W}-\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \left\langle W,H_{ij}\right\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t)},$ (20)
and $G_{v_{j},W}^{(t)}$ is the cumulative sum of the squared gradients as
$G_{v_{j},W}^{(t)}=G_{v_{j},W}^{(t-1)}+({g^{{}^{\prime}}}_{v_{j},W}^{(t)})^{2}.$ (21)
@@
This method adaptively adjusts the step size by accumulating historical gradient information. If the estimate of a variable node for a certain Pauli error changes significantly, the algorithm reduces the step size to counteract oscillation. (19) takes effect after the first iteration to ensure initial update stability. Here, $\alpha$ is the boost learning rate, providing an update impetus as the algorithm begins to accumulate gradients. $\alpha$ is relatively insensitive, so it is fixed at 5 in our simulations. While AdaGrad-BP may face issues with accumulated gradients causing the step size to approach zero after many iterations, in practice it can reduce the number of iterations, allowing most error patterns to converge before the gradients become too small.
@@
Unfortunately, combining Momentum-BP and AdaGrad-BP using Adam optimizer or other methods will not yield satisfactory results. On one hand, the Adam algorithm introduces additional hyperparameters whose optimal values for BP cannot be directly borrowed from deep learning, thereby complicating the algorithm. On the other hand, Momentum and AdaGrad approach similar message update optimizations from different perspectives, and a straightforward combination of the two can lead to overcompensation.
III-B Dynamic Initialization of Prior Probabilities
@@
BP decoding and gradient descent still have a fundamental difference, which limits the performance of the algorithms in Section III-A. Each update in gradient descent is based on the results of the previous step, whereas in BP decoding, each calculation of a posteriori probabilities accumulates the same a priori probabilities. Given the objective function as shown in (2), for simplicity using the binary domain for instance, BP calculates the LLR values as
@@
$=\Pi_{v_{j}}+\sum_{\begin{subarray}{c}c_{i^{\prime}}\in\mathcal{M}(v_{j})\backslash c_{i}\end{subarray}}m_{c_{i^{\prime}}\rightarrow v_{j}},$ (24)
where $\Pi_{v_{j}}$ is a manually set initial value of the algorithm that participates in each iteration’s a posteriori probability update. However, since the initial error probabilities are derived from a theoretical channel which is difficult to measure and estimate in experiments, the a priori probabilities can significantly affect the BP iterations.
Existing methods for altering initial probabilities include reinitialization using information obtained from previous iterations [24, 25, 33], as well as detecting oscillations during the decoding process and modifying initial values accordingly [34]. However, they either require multiple outer loops or manually set algorithm parameters. We introduce a simple but effective method for dynamically updating a priori probabilities, employing a transformed exponential weighted average, incorporating components of previous a posteriori probabilities into the a priori probabilities with a decaying factor. Hence, we call this method EWAInit-BP. EWAInit-BP under parallel scheduling is shown in Algorithm 2.
One might wonder why the first term on the RHS of (25) is each qubit’s list of initial a priori probabilities $\Pi_{v_{j}}^{(0)}$ instead of a priori probabilities from the previous iteration $\Pi_{v_{j}}^{(t-1)}$, as the latter aligns better with the definition and intuition of the exponential weighted average. In fact, by using the definition in (25) rather than the aforementioned approach, we achieve the exponentially weighted average of horizontal message updates for the a posteriori probabilities in EWAInit-BP, as stated in Theorem 1.
Algorithm 2 EWAInit-BP under Parallel Scheduling
Input: Parity-check matrix $\boldsymbol{H}$, syndrome vector $\boldsymbol{s}$, a priori LLR vectors $\{\Pi_{v_{j}}^{(0)}\}_{j=1}^{N}$, max iterations $iter_{max}$, discount factor $\alpha$.
@@
$t=1,\ \Pi^{(1)}=\Pi^{(0)},\ m_{v_{j}\to c_{i}}^{(0)}=\lambda_{H_{ij}}(\Pi_{v_{j},X}^{(0)},\Pi_{v_{j},Y}^{(0)},\Pi_{v_{j},Z}^{(0)})$
EWA Initialization when $t>1$:
for $j=1\to N$, do
$\Pi_{v_{j}}^{(t)}=\alpha\Pi_{v_{j}}^{(0)}+(1-\alpha)Q_{v_{j}}^{(t-1)}$ (25)
@@
for $j=1\to N$,
Calculate $m_{v_{j}\rightarrow c_{i}}^{(t)}$ as (8), replacing $\Pi_{v_{j},W}$ with $\Pi_{v_{j},W}^{(t)}$
@@
for $j=1\to N$, $W\in\{X,Y,Z\}$, do
Calculate $Q_{v_{j},W}^{(t)}$ as (9), replacing $\Pi_{v_{j},W}$ with $\Pi_{v_{j},W}^{(t)}$
if $Q_{v_{j},W}^{(t)}>0$ for all $e$, let $\hat{E}_{i}=I$
else, let $\hat{E}_{i}=\underset{W\in\{X,Y,Z\}}{\arg\min}Q_{v_{j},W}^{(t)}$
@@
else $t=t+1$, repeat from EWA Initialization.
Theorem 1
The a posteriori update of EWAInit-BP (Algorithm 2) realizes an exponentially weighted average of past horizontal message updates, specifically expressed as
$Q_{v_{j},W}^{(t)}$ $=\Pi_{v_{j},W}^{(0)}+\sum_{k=0}^{t}(1-\alpha)^{k}\sum_{\begin{subarray}{c}c_{i}\in\mathcal{M}(v_{j})\\ \langle W,H_{ij}\rangle=1\end{subarray}}m_{c_{i}\rightarrow v_{j}}^{(t-k)}.$ (26)
Proof.
This result can be derived by combining (25) and (9) and recursively expanding the update formula. The detailed process is provided in Appendix VI-B. ∎
Notably, EWAInit-BP and Momentum-BP share a similar recusive form, but the former is more concise (see Appendix VI-B). The primary distinction between them lies in the differing placements of EWA (or momentum). Moreover, by incorporating the EWA step into the initialization phase, EWAInit-BP avoids the issue of message coefficient inconsistency caused by performing posterior updates followed by vertical updates. In fact, EWAInit-BP also demonstrates better performance in simulations.
III-C Complexity
@@ IV Simulation Results
IV-A Breaking the Trapping Sets
@@
(b) AdaGrad-BP (parallel)
(c) AdaGrad-BP (serial)
Figure 2: Comparison of traditional LLR-BP4 and AdaGrad-BP under parallel scheduling and serial scheduling on the (4,0) trapping set. The improved BP algorithm break the symmetry of message updates and converge within 12 and 2 iterations separately.
We first compare the performance of improved BP on the trapping set shown in Fig. 1. In this example, qubit 1 and qubit 2 experience X errors, resulting in an error syndrome of $[1,0,1,0]$. Fig. 2 shows the comparison between the traditional LLR-BP4 and AdaGrad-BP in Section III-A under parallel and serial scheduling, with the maximum iteration set to 50 and the a priori probabilities set to 0.01. The hard decision result of the traditional BP oscillates continuously between all $X$ and all $I$, whereas improved BP algorithm show significantly reduced oscillation amplitude and converge within few iterations.
A more detailed comparison of the different algorithms proposed in this article is presented in Table. I. As described in Section III-A, we set $\alpha$ of AdaGrad-BP to 5 and $\gamma$ of Momentum-BP to 0. For Momentum-BP, EWAInit-BP, and MBP, we tested 10 evenly distributed values for alpha within the range [0, 1] and selected the one that yielded the best performance. The same parameter selection strategy is also applied in Section IV-B. MBP (equals to AMBP in this simulation) [21] and EWAInit-BP converges at the same fastest speed under serial scheduling but oscillates under parallel scheduling. BP-OTS [34] shows little dependency on scheduling method but its convergence depends on the manually set oscillation period $T$, which may vary in different error correction scenarios. Algorithms proposed in Section III-A, Momentum-BP and AdaGrad-BP, converge fast under both parallel scheduling and serial scheduling.
It is worth noting that EWAInit-BP fails to decode trapping sets under parallel scheduling. This is because the recursive message update formula for each bit remains symmetric, making it unable to resolve the issue within trapping sets. In practical error-correcting code decoding, this limitation may lead to non-convergence in very rare special cases. However, in most cases, messages originating from outside the trapping sets can assist in achieving convergence.
TABLE I: Iterations required by different algorithms on the (4,0) trapping set
Algorithm Parallel Scheduling Serial Scheduling
@@
(A)MBP [21] $\alpha=0.5,iter=2$
BP-OTS [34] $iter=T+2$ $iter=T+2$
EWAInit-BP $\alpha=0.5,iter=2$
AdaGrad-BP $iter=12$ $iter=2$
Momentum-BP $\alpha=0.5,iter=2$ $\alpha=0.6,iter=2$
IV-B Improved BP on Surface codes
In this subsection, we compare the performance of proposed BP improvements with other improvements for various surface codes. For all BP algorithms, the maximum number of iterations $iter_{max}$ across all code lengths is set to 100 to ensure linear asymptotic complexity.
IV-B1 Comparison of our BP optimization methods
Figure 3: Overall Performance of proposed BP algorithms with existing algorithms on a $[[85,1,7]]$ planar surface code. Momentum-BP, AdaGrad-BP, MBP and AMBP are evaluated under both parallel and serial scheduling. EWAInit-BP performs between serial MBP and serial AMBP and is almost identical to BP2-OSD-0 in this case.
A comparison of all methods proposed in this article with some existing high-accuarcy BP improvements on a $L=7$ planar surface code is illustrated in Fig. 3, using the same parameter selection strategy as described in Section IV-A. The two methods for optimizing message update rules, Momentum-BP and AdaGrad-BP, exhibit improvements over traditional LLR-BP4 at physical error rates lower than 10%, providing evidence for the validity of the gradient perspective of BP. The prior-probability-based method, EWAInit-BP, demonstrates improvements ranging from 1 to 3 orders of magnitude over traditional BP and effectively eliminates the error floor. Compared to other algorithms, EWAInit-BP achieves performance between serial MBP and serial AMBP while using hardware-friendly parallel scheduling, and in this case, performs nearly identically to BP2-OSD-0 [49], despite the latter having a time complexity of $O(n^{3})$.
The enhancements observed with Momentum-BP and AdaGrad-BP support the validity of the gradient perspective, providing a foundation for further research in this direction. Next, we will provide detailed comparisons between our best enhancement, EWAInit-BP and algorithms with similar time overheads (excluding post-processing and external loops). Although AMBP[21] has an asymptotic complexity of $O(N)$, its actual runtime overhead is high, as detailed in Appendix VI-C, which also includes the results for Adaptive EWA-BP.
IV-B2 Toric code
We construct the $[[2n^{2},1,n]]$ toric code based on its properties as a hypergraph product of an $[n,1]$ repetition code. The toric code features periodic boundaries, hence all its stabilizers have a weight of four, and any error syndromes on it exhibit translational symmetry [17], which is particularly conducive to BP and neural network decoding. We compare the performance of our EWAInit-BP and other existing methods under both parallel and serial scheduling as shown in Fig. 4 and Fig. 5, respectively. Note that in all simulations in this subsection, the parameter choices for different code lengths are consistent.